# Time Series Forecasting with LLMs: Understanding and Enhancing Model Capabilities

### v1

@@ 5.1 External Knowledge Enhancing Time Series Forecasting
We introduce a novel method to improve the performance and stability of large language models for time series forecasting. The core idea of this part is to use the knowledge obtained from pre-training of large language models to help predict. We will give the large language model some basic information about the current dataset such as the name of the dataset in the prompt, and this process does not involve data leakage (it does not tell the large language model about the period or predicted value, etc.) Assuming we define initial prompt as the original time sequence, the extra information defined as , the new prompt can be written as:
@@ 5.1.1 Implementation Details
We input the dataset’s external knowledge through prompts before the sequence’s input. The external knowledge of each dataset is presented in Appendix A.1. The external knowledge includes the description of dataset features and scenarios and the dataset’s data properties. The results are shown in Table 5.
@@ 5.1.2 Key Findings
In External Knowledge Enhancement, GPT-4 generally performs better than GPT-3.5 on MSE, MAE and MAPE, especially on AirPassengers, AusBeer and other datasets. Llama-2 significantly outperforms GPT-3.5 and GPT-4 in terms of MSE and MAE metrics on some datasets (e.g., Wooly, ETTh1, ETTm2), indicating that it can capture data features more accurately. R-squared values are shown on some datasets such as ETTh1, ETTm2, and Turkey Power, all models can provide relatively accurate predictions with R-squared values close to 1.
In the LLMTime Prediction, GPT-4 and Llama-2 perform relatively well on AirPassengers and AusBeer datasets with R values close to or above 0.5. GPT-3.5 achieves very high MSE on Sunspots and Wine datasets, which may be due to the complexity of the dataset or the poor adaptation of the model for these specific tasks. The Turkey Power dataset has a high R value on the time prediction task for all models, indicating a good model fit for this task.
5.2 Natural Language Paraphrasing of Sequences
We will be creating experiments related to natural language paraphrasing of sequences. This will be based on the fact that LLMs are not affected by the order of magnitude and size of digits Shah et al. (2023).
We have created a method to convert sequences into natural language paraphrases, the main process is using traditional natural language processing techniques to preprocess the series. For instance, if we have a discrete time series that represents temperature changes, we can denote it as , where . To do this, we extract a natural language description of the sequence based on two data points and a trend from to , "This is a discrete temperature time series. The temperature rises from to , and falls from to …". The string we get is our natural language paraphrasing sequence.
The sequence generated by the natural language paraphrasing of the text in is represented by . Finally, we paraphrase the sequence from to through the large language model; we let the LLM prediction of get , after Extraction and Processing, get , then the above completed our prediction task.
We load in the sequence and determine the movement of the sequence at each step. Then add "from xx increasing to xx," or "from xx decreasing to xx." Complete the process of modifying from time series to natural language. To convert a text sequence into a discrete sequence, we use the Reverse function, which converts the natural language series into a regular expression format through hard code.
@@ 5.2.1 Implementation Details
We use the LLMTime model with GPT-3.5-Turbo, GPT-4-turbo and Llama-2 as the core to forecast the built-in time series data set, the training set and test set ratio is 4:1. We will get the time series prediction results and the original series MSE, MAE, MAPE and R. The result is present in Table 4 compared with LLMTime forecasting results.
@@ 5.2.2 Key Findings
According to the experimental data obtained in the above table, we find that enhancing LLM through natural language paraphrasing has a certain improvement effect on time series analysis and prediction on some datasets. From the perspective of model performance comparison, the R value of GPT-3.5-Turbo in Natural Language Paraphrasing and LLMTime Prediction method is negative in some cases, indicating that the model performs poorly. GPT-4-Turbo performs better on most datasets compared to other magic methods, especially on Natural Language Paraphrasing methods, MSE, MAE, and MAPE usually decrease, while R improves. Llama-2 shows better performance on some datasets, such as lower MSE and MAE and better MAPE and R in the Natural Language Paraphrasing method of the AusBeer dataset.
6 Related Works
6.1 Time Series Prediction
Time series prediction is an important field in the field of artificial intelligence. The unified idea of time series prediction is to analyze past and future series trends. It is used in finance and business to predict the trend of indexes and stocks, in the field of intelligent manufacturing to predict anomaly detection and power load, and in the field of meteorology, agriculture, and navigation to predict temperature, humidity, and climate Singh and Malhotra (2023); Munir et al. (2018); Shen et al. (2020); Braei and Wagner (2020). We categorize related research within this domain into three categories: traditional time series prediction, time series prediction based on machine learning, and deep learning.
| Metrics | GPT4-MAPE | GPT4-R2 | GPT3.5-MAPE | GPT3.5-R2 | Trend Strength | Seasonal Strength |
| GPT4-MAPE | ||||||
| GPT4-R2 | ||||||
| GPT3.5-MAPE | ||||||
| GPT3.5-R2 | ||||||
| Trend Strength | ||||||
| Seasonal Strength |
Two commonly used methods for traditional time series analysis are the ARIMA method Box and Pierce (1970) and the exponential smoothing method Gardner Jr (2006). The ARIMA model is a classic forecasting method that breaks down a time series into auto-regressive (AR), difference (I), and moving average (MA) components to make predictions. On the other hand, exponential smoothing is a straightforward yet effective technique that forecasts future values by taking a weighted average of past observations. The ARIMA model requires testing the stationarity of data and selecting the right order. On the other hand, the exponential smoothing method is not affected by outliers, it is only suitable for stationary time series, and its accuracy in predicting future values is lower compared to the ARIMA model.
| Dataset | Period | Real | Median | |||||||||
| AirPassengersDataset | 1 | 12 | 18 | 21 | 7 | 6 | 18 | 18 | 4 | 6 | 12 | 12 |
| AusBeerDataset | 24 | 10 | 6 | 6 | 1 | 3 | 6 | 7 | 3 | 21 | 4 | 6 |
| GasRateCO2Dataset | 7 | 36 | 24 | 72 | 28 | 34 | 20 | 7 | 16 | 14 | 39 | 20 |
| MonthlyMilkDataset | 6 | 6 | 25 | 11 | 6 | 2 | 7 | 34 | 8 | 7 | 12 | 7 |
| SunspotsDataset | 1 | 12 | 18 | 21 | 7 | 6 | 18 | 18 | 4 | 6 | 12 | 12 |
| WineDataset | 1 | 11 | 12 | 11 | 12 | 7 | 6 | 8 | 10 | 12 | 12 | 10 |
| WoolyDataset | 10 | 5 | 6 | 7 | 16 | 11 | 5 | 11 | 3 | 4 | 4 | 5 |
| HeartRateDataset | 13 | 16 | 6 | 12 | 13 | 11 | 12 | 32 | 32 | 64 | 144 | 12 |
Time series analysis based on linear regression, decision tree random forest, and other methods can make predictions for future series through the analysis and learning of existing series, as more complex time series prediction methods based on deep learning, RNN, CNN, and LSTM have played a huge role in many applications. Especially in the field of sequence prediction, many variants of LSTM have emerged, such as bidirectional LSTM Graves et al. (2005), gated recurrent unit Chung et al. (2014), long short-term memory network Malhotra et al. (2015), and deep recurrent neural network Pascanu et al. (2013). This method improves prediction accuracy but takes longer to train and is less prone to overfitting than the original LSTM.
6.2 LLMs for Time Series Prediction
Given the notable achievements of LLMs across various domains, it is prudent to incorporate time series prediction tasks with LLMs. Similar to (Sun et al., 2023), we categorize contemporary research within this domain into two categories: LLM-for-Time Series and Time Series-for-LLM.
LLM-for-Time Series involves either pre-training a foundational large language model or fine-tuning existing LLMs by leveraging extensive time-series data Rasul et al. (2023); Garza and Mergenthaler-Canseco (2023); Das et al. (2023); Cao et al. (2023). For instance, (Rasul et al., 2023) aimed to build the foundational models for time series and investigate its scaling behavior. Chang et al. (2024) proposed a two-stage fine-tuning strategy for handling multivariate time-series forecasting. Although these studies contribute significantly to understanding foundational models, they demand considerable computing resources and expertise in fine-tuning procedures. Moreover, the details of the model may not be disclosed for commercial purposes Garza and Mergenthaler-Canseco (2023), which impedes future research. Additionally, in scenarios with limited data available, there is insufficient information for training or fine-tuning. Therefore, in this paper, we concentrate on the Time Series-foR-LLM, harnessing fixed LLMs for the time-series forecasting task.
@@ 6.2 LLMs for Time Series Prediction
| Wine | 11403.89 | 96.95 | 37.04 | 30488.60 | 388.28 | 15.83 | |
| Wooly | 12110.16 | 33.23 | 4.07 | 526903.08 | 574.58 | 12.00 | |
| HeartRate | 4.38 | 0.55 | 0.57 | 76.83 | 7.15 | 7.42 | |
@@
| Wine | 56138.87 | 54.67 | 23.63 | 22488.17 | 253.08 | 9.98 | |
| Wooly | 18063.64 | 11.06 | 25.06 | 942987.19 | 871.64 | 18.55 | |
| HeartRate | 11.64 | 1.21 | 1.30 | 988.14 | 26.57 | 29.22 | |
| Llama-2 | AirPassengers | 751.34 | 6.77 | 1.53 | 1286.25 | 28.04 | 6.07 |
@@
| Wine | 102434.52 | 852.97 | 34.72 | 951194.94 | 240.08 | 9.45 | |
| Wooly | 12180.05 | 83.99 | 16.92 | 675062.52 | 736.04 | 15.83 | |
| HeartRate | 49.8 | 5.84 | 6.53 | 75.58 | 7.11 | 7.94 | |
@@
Time Series-for-LLM centres around using existing LLMs and designing mechanisms, such as crafting appropriate prompts or reprogramming inputs, to handle time series data effectively (Gruver et al., 2023; Sun et al., 2023; Jin et al., 2023; Xue and Salim, 2023). Sun et al. (2023) tokenized the time series and managed to embed those tokens, and Jin et al. (2023) reprogrammed the time series data with text prototypes before feeding them to the LLMs. These studies illuminated the characteristics of time series data and devised methods to align them with LLMs. However, they lack an analysis of the ability and bias in forecasting time series. The most related work to us is Gruver et al. (2023), though, it lacks a quantitative analysis of the preference for the time series in LLMs, and it failed to explore the impact of input forms and prompt contents, such as converting the numerical time series into the natural language sequences and incorporating the background information into the prompt. Our work fills the gap, and we expect our work to be the benchmark for time-series analysis and provide insights for subsequent research.
7 Conclusions
@@
Our research has unearthed key preferences of large language models (LLMs) in the domain of time series forecasting, revealing a proclivity for data with distinct trends and seasonal patterns. Through a blend of real and synthetic datasets, coupled with counterfactual experiments, we have demonstrated LLMs’ improved forecasting performance with time series that exhibit clear periodicity. Our proposed strategies incorporating external knowledge and transforming numbers into natural language have both enhanced LLMs’ predictive accuracy. Despite encountering challenges with multi-period time series, our work provides a pivotal foundation for future advancements, offering actionable insights and methods to refine the application of LLMs in time series analysis.
@@
This study may be limited in the following ways. First, limitations in the scope of the dataset and large language models may not capture the full variability of the results with a wider array. In addition, some experimental sessions lack a comparison with hard-coded solutions, and there is a gap in understanding the performance of LLMs compared to traditional programming methods. Furthermore, the inability to categorize datasets by type and conduct specific types of experiments limits insight into the performance of the model in different data domains. These limitations suggest that the results could benefit from more extensive experiments and more nuanced analyses, which underscores the need to expand future research.
@@ Appendix A Appendix
A.1 The External Knowledge incorporated in the Prompt
@@
(2) AusBeerDataset: This is a quarterly series of beer production, with each value representing the kilolitres of beer produced in that quarter.
@@
(4) MonthlyMilkDataset: This is a time-series data set describing monthly milk production. Each is the average number of tons of milk each cow produces during the month.
(5) SunspotsDataset: This is a dataset that records the number of sunspots in each month, where each data is the number of sunspots in that month.
@@
(9) Weather: This dataset contains weather-related data such as temperature, precipitation, humidity, etc. It is commonly used in weather forecasting and climate research.
(10) COVID Deaths: This is a collection of data documenting related deaths during the COVID-19 pandemic, including information on the number of deaths, infection rates, etc., for use in outbreak analysis and health policy development.
(11) Solar Weekly: This dataset contains weekly data on solar power generation, which can be used to analyze and forecast renewable energy production.
(12) Tourism Monthly: This dataset records monthly tourism-related data such as the number of tourists, tourism revenue, etc., for tourism trend analysis and prediction.
(13) Australian Electricity Demand: This deals with electricity demand data in Australia, which can be used for electricity market analysis and grid management.
(14) Pedestrian Counts: Pedestrian count data are recorded for specific areas, such as city streets, for urban planning and traffic management.
(15) Traffic Hourly: Provides hourly traffic flow data for traffic forecasting and urban planning.
(16) Hospital: This dataset may contain hospital-related data such as admission rates, bed occupancy rates, etc., for medical resource planning and public health research.
(17) Fred MD: It is probably a collection of economic data, such as gross domestic product (GDP), unemployment rate, etc., used for macroeconomic analysis and forecasting.
(18) Tourism Yearly: Similar to "Tourism Monthly", but provides annual tourism data for longer-term analysis of tourism trends.
(19) Tourism Quarterly: Provides quarterly tourism data between monthly and annual data for medium-term tourism trend analysis.
(20) US Births: A dataset that records birth rates in the United States, including information such as the number of births, for use in population research and social policy making.
(21) NN5 Weekly: This may be a weekly financial or economic dataset, including stock market data, economic indicators, and so on.
(22) Traffic Weekly: Provides weekly traffic flow data for long-term traffic pattern analysis and planning.
(23) Saugeenday: This may be a daily degree dataset for a specific region, such as the Saugeen region, and the data involved may include environmental, meteorological, or socioeconomic data.
(24) CIF 2016: This may be a specific economic or financial time series dataset collected in 2016 for economic forecasting and analysis.
(25) Bitcoin: This dataset deals with the price and transaction data of Bitcoin and is used for cryptocurrency market analysis and forecasting.
(26) Sunspot: Data recording sunspot activity for astronomical research and prediction of the solar cycle.
(27) NN5 Daily: This is similar to NN5 Weekly, but provides daily data that may refer to financial markets or economic indicators.
A.2 Details of the experiments on the LLMs’ Preferences
In this subsection, we provide a comprehensive overview of the experiments conducted to investigate the preferences of LLMs for input time series data. We first describe both the real and synthesized datasets we use, and then detail the methods we use to investigate the preferences of LLMs.

### v3

@@ 5.1 External Knowledge Enhancing Time Series Forecasting
We introduce a novel method to improve the performance of large language models for time series forecasting. The core idea of this part is to use the knowledge obtained from the pre-training stage to help predict. We provide the large language model with some basic information about the current dataset such as the background of the data collection, and this process does not involve data leakage. We incorporate our tests on the data leakage in Appendix A.4. It is noted that we do not provide the LLMs with any statistical information such as the periods or trends. This approach ensures that the LLMs forecast the time series entirely based on the data and their prior knowledge. Let denote the initial prompt representing the original time sequence, and let denote the additional information. Consequently, the new prompt can be expressed as: .
@@ 5.1.1 Implementation Details
We input the dataset’s external knowledge through prompts before the sequence’s input. The external knowledge of each dataset is presented in subsection A.1. The results are shown in Table 8, where LLMTime Prediction refers to the approach described by [10] without any modifications.
@@ 5.1.2 Key Findings
As shown in Table 8, this method achieves improved performance in most scenarios. Besides, GPT-4-turbo generally performs better than GPT-3.5-turbo on MSE, MAE, and MAPE, especially on AirPassengers, AusBeer, and other datasets. Llama-2 significantly outperforms GPT-3.5-turbo and GPT-4-turbo in terms of MSE and MAE metrics on some datasets (e.g., Wooly, ETTh1, ETTm2), indicating that it can capture data features more accurately. Using External Knowledge Enhancing, Gemini outperforms other models on MonthlyMilk, Sunspots, Wooly, and HeartRate Datasets, but performs poorly on other datasets.
5.2 Natural Language Paraphrasing
In this subsection, we conduct experiments on the natural language paraphrasing of the input time sequences. This strategy capitalizes on the advanced abilities of large language models in handling language sequences. It is motivated by the fact that LLMs are insensitive by the order of magnitude and size of digits [18].
We use natural language to describe the trend between consecutive values. For instance, given a time series where , we describe the trend from to as follows: "The value rises from to , and falls from to …". The string we get here is our natural language paraphrasing sequence. After generating responses based on the string, we extract the values from the text and construct the predicted time series.
| Model | Dataset | Period | Real | Mode | |||||||||
| AirPassengersDataset | 24 | 24 | 7 | 24 | 12 | 24 | 11 | 24 | 24 | 24 | 12 | 24 | |
| WineDataset | 11 | 12 | 24 | 24 | 24 | 20 | 24 | 24 | 24 | 24 | 12 | 24 | |
| GPT-3.5-turbo | MonthlyMilkDataset | 6 | 9 | 12 | 9 | 12 | 12 | 12 | 12 | 12 | 11 | 12 | 12 |
| WoolyDataset | 4 | 3 | 4 | 3 | 3 | 4 | 3 | 3 | 6 | 3 | 4 | 3 | |
| AusBeerDataset | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 3 | 4 | 3 | |
| AirPassengersDataset | 11 | 12 | 12 | 4 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | |
| WineDataset | 10 | 12 | 24 | 12 | 6 | 12 | 12 | 24 | 12 | 12 | 12 | 12 | |
| Gemini-Pro-1.0 | MonthlyMilkDataset | 16 | 12 | 12 | 12 | 12 | 39 | 12 | 11 | 12 | 12 | 12 | 12 |
| WoolyDataset | 5 | 7 | 4 | 5 | 4 | 4 | 4 | 4 | 5 | 6 | 4 | 4 | |
| AusBeerDataset | 4 | 4 | 4 | 2 | 5 | 5 | 4 | 3 | 5 | 7 | 4 | 4 | |
| AirPassengersDataset | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | 12 | |
| WineDataset | 7 | 6 | 6 | 6 | 7 | 7 | 6 | 6 | 6 | 6 | 12 | 6 | |
| GPT-4-turbo | MonthlyMilkDataset | 10 | 12 | 12 | 12 | 12 | 14 | 12 | 12 | 12 | 12 | 12 | 12 |
| WoolyDataset | 5 | 5 | 7 | 5 | 5 | 5 | 7 | 5 | 5 | 4 | 4 | 5 | |
| AusBeerDataset | 4 | 4 | 4 | 4 | 6 | 4 | 4 | 4 | 6 | 4 | 4 | 4 | |
@@ 5.2.2 Key Findings
According to the results in Table 3, we find that enhancing LLM through natural language paraphrasing improves time series forecasting on most datasets. For instance, GPT-3.5-turbo and GPT-4-turbo perform better on most datasets, especially on Natural Language Paraphrasing methods. Gemini outperforms other LLMs on Wooly and Ausbeer datasets but underperforms on others with natural language paraphrasing. All these results demonstrate the superior performance of our methods.
@@
| (GPT-3.5-turbo-1106) | HeartRate | 4.38 | 0.55 | 0.57 | 76.83 | 7.15 | 7.42 |
| Istanbul-Traffic | 224.17 | 3.74 | 8.81 | 335.05 | 6.75 | 11.68 | |
| ETTh1 | 1.21 | 0.48 | 54.17 | 5.64 | 2.71 | 1.625 | |
| ETTm2 | 0.81 | 0.36 | 27.33 | 3.46 | 2.17 | 1.178 | |
@@
| (GPT-4-turbo-preview) | HeartRate | 11.64 | 1.21 | 1.30 | 988.14 | 26.57 | 29.22 |
| Istanbul-Traffic | 176.91 | 3.88 | 9.67 | 195.33 | 5.53 | 10.03 | |
| ETTh1 | 1.20 | 0.49 | 47.62 | 4.73 | 1.53 | 3.282 | |
| ETTm2 | 0.45 | 0.27 | 23.62 | 2.30 | 1.034 | 1.607 | |
| Llama-2 | AirPassengers | 751.34 | 6.77 | 1.53 | 1317.9 | 55.49 | 11.18 |
@@
| (llama-2-13B) | HeartRate | 49.8 | 5.84 | 6.53 | 75.58 | 7.11 | 7.94 |
| Istanbul-Traffic | 306.80 | 5.39 | 7.24 | 438.28 | 7.28 | 9.81 | |
| ETTh1 | 1.47 | 0.87 | 58.34 | 4.84 | 1.79 | 3.178 | |
| ETTm2 | 0.84 | 0.41 | 29.86 | 3.31 | 2.07 | 2.153 | |
| Gemini-Pro-1.0 | AirPassengers | 4474.54 | 31.54 | 7.02 | 6392.21 | 63.57 | 14.03 |
| AusBeer | 278.45 | 10.05 | 2.29 | 397.78 | 14.36 | 3.27 | |
| GasRateCO2 | 13.29 | 2.50 | 4.38 | 18.99 | 3.57 | 6.46 | |
| MonthlyMilk | 440.29 | 11.91 | 1.39 | 628.98 | 17.01 | 1.99 | |
| Sunspots | 438.29 | 10.47 | 1.21 | 626.03 | 14.94 | 1.73 | |
| (gemini-1.0-pro) | HeartRate | 40.57 | 4.20 | 4.67 | 57.96 | 6.01 | 6.66 |
| Istanbul-Traffic | 267.43 | 5.69 | 8.37 | 321.56 | 7.32 | 9.71 | |
| ETTh1 | 1.17 | 0.74 | 54.86 | 4.84 | 1.79 | 3.178 | |
| ETTm2 | 0.88 | 0.39 | 21.82 | 3.31 | 2.07 | 2.153 | |
6 Related Work
In this section, we review two lines of research that are most relevant to ours.
6.1 Traditional Time Series Forecasting
Two commonly used methods for traditional time series analysis are the ARIMA method [2] and the exponential smoothing method [7]. The ARIMA model is a classic forecasting method that breaks down a time series into auto-regressive (AR), difference (I), and moving average (MA) components to make predictions. On the other hand, exponential smoothing is a straightforward yet effective technique that forecasts future values by taking a weighted average of past observations. The ARIMA model requires testing data stationarity and selecting the right order. However, the exponential smoothing method is not affected by outliers, it is only suitable for stationary time series, and its accuracy in predicting future values is lower than the ARIMA model.
6.2 LLMs for Time Series Forecasting
The first family of methods involve either pre-training a foundational large language model or fine-tuning existing LLMs by leveraging extensive time-series data [17, 8, 6, 3]. For instance, [17] aimed to build the foundational models for time series and investigate its scaling behavior. [4] proposed a two-stage fine-tuning strategy for handling multivariate time-series forecasting. Although these studies contribute significantly to understanding foundational models, they require considerable computing resources and expertise in fine-tuning procedures. Moreover, the details of the model may not be disclosed for commercial purposes [8], which impedes future research. Additionally, in scenarios with limited data available, there is insufficient information for training or fine-tuning.
@@ 6.2 LLMs for Time Series Forecasting
In contrast, the second family of methods does not involve model parameter finetuning. These methods either create appropriate prompts or reprogramme inputs, to effectively handle time series data [10, 19, 14, 21]. [19] tokenizes the time series and manages to embed those tokens, and [14] reprogrammed the time series data with text prototypes before feeding them to the LLMs. These studies illuminate the characteristics of time series data and devise methods to align them with LLMs. However, they lack an analysis of the ability and bias in forecasting time series. The most related work to us is [10], though it lacks a quantitative analysis of the preference for the time series in LLMs, and it fails to explore the impact of input forms and prompt contents, such as converting the numerical time series into the natural language sequences and incorporating the background information into the prompt. Our work fills the gap, and we expect our work to be the benchmark for time-series analysis and provide insights for subsequent research.
7 Conclusions and Future Work
@@
In this work, we investigate the key preferences of LLMs in the domain of time series forecasting under the zero shot setting, revealing a proclivity for data with distinct trends and seasonal patterns. Through a blend of real and synthetic datasets, coupled with counterfactual experiments, we have demonstrated LLMs’ improved forecasting performance with time series that exhibit clear periodicity. Besides, our results indicate that LLMs struggle with multi-period time series datasets, as they face difficulty in recognizing the distinct periods within them. Our findings also suggest that large language models are more sensitive to the segment of input sequences closer to the last known data than other locations. Lastly, experimental results indicate that our proposed strategies of incorporating external knowledge and transforming numerical sequences into natural language formats have yielded substantial improvements in accuracy.
@@
This study may be limited in the following ways. First, limitations in the scope of the dataset and large language models may not capture the full variability of the results with a wider array. In addition, some experimental sessions lack a comparison with hard-coded solutions, and there is a gap in understanding the performance of LLMs compared to traditional programming methods. Furthermore, the inability to categorize datasets by type and conduct specific types of experiments limits insight into the model’s performance in different data domains. These limitations suggest that the results could benefit from more extensive experiments and more nuanced analyses, underscoring the need to expand future research.
@@ Appendix A Appendix
A.1 Dataset description and the External Knowledge incorporated in the Prompts
In this subsection, we briefly introduce the datasets we use, which also serve as the external knowledge incorporated into the prompts. Following [10], we downsample the input series to an hourly frequency, yielding a total of 267 observations and resulting in relatively small datasets. Additionally, we incorporate Memorization datasets published after September 2021, the cutoff date for GPT-3.5-turbo, to demonstrate the effectiveness of TimeLLM and our proposed methods. Finally, we implemented univariate time series forecasting to predict the ’OT’ feature on the ETTh1 and ETTm2 datasets, focusing on the last 96 steps of the test set.
A.1.1 Darts [11]
@@
(2)AusBeerDataset: This is a quarterly series of beer production, with each value representing the kiloliters of beer produced in that quarter.
@@
(4)MonthlyMilkDataset: This time-series data set describing monthly milk production. Each is the average number of tons of milk each cow produces during the month.
(5)SunspotsDataset: This dataset records the number of sunspots each month, where each data is the number of sunspots in that month.
@@
(9)ETTh1: This is a time series dataset containing high-frequency energy data of a certain region in China, which is mainly used for energy load forecasting and related time series analysis research.
(10)ETTm2: Similar to ETTh1, ETTm2 is also an energy time series dataset, but the data frequency or region covered may be different, which is also used for the analysis and prediction of energy consumption.
A.1.2 Memorization Datasets [10]
(11)TSMCStockDataset: This is historical trading data about Taiwan Semiconductor Manufacturing Corporation (TSMC) stock, containing information such as share price, volume, and date, and is commonly used in financial analysis and stock market forecasting research.
(12)TurkeyPowerDataset: This is a time series dataset on national electricity consumption in Turkey, which records the electricity usage in Turkey over a period of time and is often used to analyze and forecast electricity demand.
(13)IstanbulTrafficDataset: This dataset offers hourly Traffic Index data for Istanbul from October 2022 to May 2023.
A.1.3 Monash Datasets [9]
(14)US Births Dataset: This dataset contains the number of births in the US from 1969 to 1988.
(15)Saugeen River Flow Dataset: This dataset contains the daily mean flow of the Saugeen River at Walkerton in cubic meters per second from 1915 to 1979.
A.2 Implementation Details
In this subsection, we provide a comprehensive overview of the experiments conducted to investigate the preferences of LLMs for input time series data. We first describe both the real and synthesized datasets we use and then detail the methods we use to investigate the preferences of LLMs.